OpenStudy (anonymous):

Will give medle Solve. –5t ≥ 65 A. t ≥ 13 B. t ≤ 13 C. t ≥ –13 D. t ≤ –13

ganeshie8 (ganeshie8):

\(\large –5t\ge 65\)

ganeshie8 (ganeshie8):

divide both sides by "-5"

ganeshie8 (ganeshie8):

wat do u get ? :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

1 and 13

ganeshie8 (ganeshie8):

\(\large –5t\ge 65\) /-5 /-5

OpenStudy (anonymous):

so b

ganeshie8 (ganeshie8):

yes ! \(\large t \color{red}{\le} -13\)

ganeshie8 (ganeshie8):

you're almost right, just pay attention to signs also !

ganeshie8 (ganeshie8):

65/-5 is -13 okay ?

OpenStudy (anonymous):

ok so it is B right?????

ganeshie8 (ganeshie8):

answer is D

OpenStudy (anonymous):

o ok

ganeshie8 (ganeshie8):

good try though :)

OpenStudy (isaiah.feynman):

@Limpers Any time you solve an inequality and you DIVIDE OR MULTIPLY BY a negative number, you reverse the sign. A simple example...